Prove that the map f : C → R


2


, x + iy 7→ (x, y) is a bijection, i.e., a cardinal equivalence


C ≈ R


1
Expert's answer
2021-11-09T11:59:39-0500

Let us prove that the map "f : \mathbb C → \R^2,\ f(x + iy)= (x, y)," is a bijection.

Let "f(x_1 + iy_1)=f(x_2 + iy_2)." Then "(x_1, y_1)=(x_2, y_2)," and hence "x_1=x_2" and "y_1=y_2." We conclude that "x_1 + iy_1=x_2 + iy_2," and thus the map "f" is injective.

Let "(a,b)\in\R^2" be arbitrary. Then for "z=a+ib" we get that "f(z)=f(a+ib)=(a,b)," and consequently, "f" is surjective.

We conclude that "f : \mathbb C → \R^2" is a bijection.